OpenStudy (anonymous):

You are a student who plans to attend the amusement park for two days. Now figure out the cost. You won a coupon for 50% off a 1-day pass but it cannot be combined with other offers. The amusement park also offers a $5 student discount; however, you cannot get the $5 student discount off the 2-day pass. Option 1: Buying a ticket at 50% off today and then purchasing a new 1-day ticket at the student price tomorrow. Option 2: Buying a 2-day pass and not using the coupon at all. 1 day at the park for $44.95 2 days for $71.95 3 days for $94.95 which one is cheaper ? show work pleas

OpenStudy (anonymous):

option 1 44.95-50%= 22.475 and then 44.95-5%=42.7025 sum 65.1775

OpenStudy (anonymous):

wait now let me solve option 2 and see which one is cheaper hold on

OpenStudy (anonymous):

okay so option 2 will obviously be71.95, but for option one, i made an error it 5$ off not 5%, so it would be 44.95-5$=39.95, sum will be 62.425 which still makes option 1 cheaper